• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name Ulmer

    Log entry time 13:20:21 on November 1,2001

    Entry number 75567

    This entry is a followup to: 75560

    keyword=More on OTR

    Arun Saha pointed out the earlier entry of J. LeRose (Entry 75501) which says that the OTR's are currently non-functional due to problems with the filters. Nonetheless, the OTR's are currently in the beam path, but Arun has reassured me that their effect on the beam is minimal enough to warrant keeping them in.
    See also the entry of Simon Sirca (Entry 75563).